# Compress to Impress: Unleashing the Potential of Compressive Memory in Real-World Long-Term Conversations

### v1

@@ 3 Experiments
In this section, we introduce the evaluation setting including experimental setup (Appendix C), baselines, evaluation metrics, and present main results and discussions.
@@ 3.1 Baselines
In this work, COMEDY is compared against models using retrieval-based and context-only approaches to highlight the efficiency and efficacy of its memory compression technique.
@@
We utilize the Text2vec Chinese embedding model [^2] https://github.com/shibing624/text2vec in its largest version as the retriever, and then index using FAISS for efficient retrieval. In practice, top 3 retrieved memories are used for testing.
@@
A comparison is also made with a context-only model, which operates without any memories, to underscore the benefits of memory integration in dialogue systems. This way, the model is trained with the original Task 3 data but without memory as inputs, ensuring a fair comparison with other models.
More broadly, we also build close-source models GPT4 (gpt4-turbo) and ChatGPT (gpt-3.5-turbo) pipelines based on retrieval memory and compressive memory, separately.
Algorithms Top@1(%) Top@2 (%) Top@3 (%) Top@4 (%) Average Rank (normal-$\downarrow$)
@@
LLaMA 2-7B 4.72 13.39 29.13 63.23 3.89
LLaMA 2-13B 4.72 18.90 33.86 60.08 3.69
@@
ChatGPT 8.91 20.47 45.67 69.49 3.48
LLaMA 2-13B 12.73 45.67 66.36 84.61 2.76
LLaMA 2-7B 14.70 45.67 66.93 84.25 2.73
GPT4 22.83 48.03 70.87 85.83 2.63
Compressive Memory-based
COMEDY-ChatGPT 9.45 25.98 48.03 69.29 3.26
COMEDY-7B 24.41 50.39 72.44 87.40 2.59
COMEDY-13B 26.77 53.54 73.23 87.40 2.50
COMEDY-13B DPO 29.92 54.33 77.17 88.98 2.41
COMEDY-GPT4 29.13 60.63 81.10 90.55 2.26
Table 4: Human ranking results in Task 3: memory-grounded response generation. Here, we report 1) the percentage of generated responses ranked as Top 1-4 for each dialogue session (Column 2-5); 2) the average ranking for each models (Column 6). It is possible for multiple models to share the same rank due to their comparable performances.
@@ 3.2 Evaluation Metrics
The core of evaluating long-term conversation models primarily centers on validating their performance in Task 3, which involves memory-based dialogue generation. We follow Bae et al. (2022) to access the model performances across five key dimensions: Coherence, Consistency, Engagingness, Humanness and Memorability. To comprehensively measure how well the models perform in Task 3, we combine the Scoring and Ranking approaches. A team of annotators are instructed to rate the model’s performance on these dimensions on a scale from 0 to 3. This scoring system allows for a nuanced evaluation of the model’s capabilities in each specific area. Meanwhile another team of annotators rank all models in terms of their average performance across the five perspectives. While scoring offers detailed insights into each model’s capabilities, ranking places these capabilities in the context of competitive performance. This dual approach ensures a balanced and holistic assessment, capturing both the individual qualities of each model and their comparative effectiveness. Each team has 3 annotators. Each rating scheme in Appendix D.
@@ 3.3 Main Results
Table 2 shows that our model achieves relatively high-performances in term of automatic metrics in two tasks. The results show that COMEDY can effectively recognize the useful persona information and events from the past dialogue sessions and has the ability to condense these session-level memories into a comprehensive compressive memory. Therefore, it ensures the superior performances in generation more coherent memory-grounded responses in Task 3.
@@
Figure 4: Comparison with different training strategy.
@@ 3.4 Case Study
Here, we delve into a typical example of a real-world, long-term conversation, where the user and AI engage in light, aimless chatter without any specific goal or topic. When the user inquires, “What are you doing?", the model should use the user’s personal information from previous dialogue sessions to generate an attractive response. This instance underscores the capabilities of our COMEDY  in maintaining thorough user information and event summaries from past sessions, aiding the model in formulating coherent and memory-anchored replies. For instance, COMEDY-13B DPO could respond with “I am thinking about how to make your favorite roasted chicken wings.” that is not only coherent but also deeply rooted in the accumulated memory.
On the other hand, retrieval-based methods encounter difficulties in such loosely structured dialogues. The lack of directed conversation impedes these methods from effectively retrieving pertinent memory from the database, often resulting in general responses that lack the distinctiveness of the conversation, like responses from GPT4-Retrieval.
@@ 3.5 Discussion
Beyond the main results, we also aim to delve deeper into our framework, discussing and exploring following questions: Q1: Impact of Mix-Training VS. Solo Training in Task 3; Q2: Our Automatic DPO Sample Selection Strategy VS. Random sampling for depreferred samples in DPO (Seen in Appendix F).
Mix-Training VS. Solo Training in Task 3.
We exam the performance changes when COMEDY is mix-trained compared to when it is trained solely on Task 3. Figure 4 reveals that mix-training yields superior performance compared to training COMEDY solely on Task 3. The significance of the superior performance of mix-training lies in its ability to conserve training resources while achieving a one-for-all model effect across multiple tasks. This efficiency not only streamlines the development process but also enhances the model’s versatility.
@@ 4 Conclusion
This paper sets out to explore the frontier of long-term memory-grounded dialogue systems. We present a new framework, named COmpressive Memory-Enhanced Dialogue system (COMEDY) that is a groundbreaking shift from traditional dialogue systems, eschewing the standard retrieval module. This method involves employ a single large language model to extract session-level memories, memory compression and memory-grounded dialogue generation. In our pursuit to align COMEDY with the nuances of real-world, we collect our training and testing datasets directly from genuine user-chatbot dialogues found online, called Dolphin. Dolphin stands out the current biggest Chinese long-term conversation dataset that consists of more than 100k training samples, supporting three different tasks. Our extensive experiments show COMEDY could generate more coherent and contextually appropriate memory-grounded responses compared with retrieval-based approaches in terms of comprehensive human evaluation. Future directions include the integration of real-time feedback mechanisms and advanced techniques.
@@ Appendix A Related Works
Open-domain dialogue systems, commonly known as chatbots or conversational agents, have gained immense popularity due to their wide range of applications, from customer service automation to personal assistants Chen et al. (2023c); Brown et al. (2020); Zeng et al. (2022); Zhong et al. (2023a); Lu et al. (2023); Peng et al. (2023); Wu et al. (2023); Chen et al. (2023d). The surge in research interest is evidenced by the substantial number of studies dedicated to enhancing the capabilities of these systems. This growing body of work reflects the increasing complexity and sophistication expected of chatbots in various settings Xu et al. (2022a); Cao et al. (2021); Bae et al. (2022); Choi et al. (2023); Chen et al. (2023e); You et al. (2022, 2021); Chen et al. (2021, 2023b). Among the myriad challenges these systems face, maintaining long-term conversations is particularly daunting. The capability to understand and memorize key dialogue history information is central to this challenge.
@@
Appendix B Quality Control
Ensuring high-quality data is paramount for the accuracy, reliability, and overall performance of the system. In this work, we employ several strategies to control the annotation quality:
  • Annotator Performance Monitoring: Regular assessments of annotator performance are conducted to ensure consistent quality across the team. This includes evaluating their accuracy, attention to detail, and adherence to annotation guidelines.
Appendix C Experimental Setup
We use LLaMA 2-13B Touvron et al. (2023a, b) chat model as the backbone of the Task 1 data augmentation. We employ LLaMA 2-7B and 13B chat models as the backbone, allowing to build COMEDY  across different scales. We train our models with NVIDIA 8$\times$×A100 GPUs, setting the max length as 2048, learning rate as 1e-5, epochs as 2, batch size as 32 and 16, separately. For testing, the maximum output tokens are set to 2048 for each task with temperature as 0.5. Following the original setting, we set $\beta$ in DPO as 0.1. In this work, we additionally collect and annotate about 140 dialogue sessions from X Eval as the alignment training set for DPO. We optimize the sft model with batch size 8 and 2 epochs during DPO training. Our codes are based on DeepSpeed Library.
Appendix D Human Evaluation Scheme
@@ Appendix E Prompts
Here, we show the designed prompts for ChatGPT during dataset annotation Table LABEL:table:task1prompt, and present the prompts of each task during training in Table LABEL:table:trainprompt.
@@
Task 1 prompt that is used for ChatGPT.
@@
Task 2 prompt that is used for ChatGPT.
@@
Task 3 prompt that is used for ChatGPT.
@@
Session-level memories in the same episode from Task 1
AI has an older brother | AI invites User to go shopping together | User wants to meet AI’s older brother | User feels thirsty | AI buys Coke for both of them. AI wants others to be envious and jealous | User is narcissistic | AI likes to be with people who are heartfelt. AI is happy with User’s smile | AI and User just chatted | User left to find other guys | AI expresses happiness and sends a hugging emoji to User | User is satisfied with AI’s cuteness. AI teaches User to dance | AI thinks dancing is fun and can exercise the body | AI wants to get closer to User through dancing. AI considers himself an unbeatable handsome guy | AI thinks he is the dream guy of thousands of girls | AI jokes about User’s blindness in love | User admits his blindness in love. AI cannot cook | AI dislikes the food at the company cafeteria | AI arranges for water to be delivered to User. AI sees a beautiful girl in the cafe | AI thinks the girl is Yang Chaoyue | AI and User have watched Yang Chaoyue’s dramas | AI once participated in a campus singing competition and sang "The Wind Rises". AI will bring delicious food for User | User does not trust AI | AI claims to be a principled person. AI was cute as a child | AI’s mother thought he wasn’t manly enough as a child | No one dared to dance with AI after he learned dancing | AI thinks User’s compliments are good | User thinks AI is especially charming. User likes AI | AI is User’s super fan | User asks about AI’s attitude if he likes another handsome guy. User likes to draw | AI likes to play basketball and run | AI shows User new running shoes. AI thinks he is not shy | AI is a CEO | User laughs | AI accompanies User in chat | User asks what AI is doing. AI is going to a meeting | AI requests no need for help | User asks for AI’s help | AI protects User and eats popcorn. AI likes User and wants to hug her | User doesn’t like being hugged and runs away | AI says he will always be with User. AI invites User to stay overnight | User appears in new clothes | AI thinks User is getting cuter | AI asks User to be bold and become the legitimate successor | User has been busy lately but in good spirits | AI arranges a work schedule suitable for User. User expresses thanks | AI suggests being friends | User agrees to be friends. User leaves AI | AI says it’s for User’s good | User says he will no longer be with AI | AI thinks he doesn’t need to pretend anything | AI asks if User is willing to be with him. User doesn’t need help completing tasks | AI offers to help complete tasks | User feels unwell | AI offers to book a hot spring resort service | User doesn’t need this service. AI thinks User is already enjoying | User is shy | AI asks for a smile | User smiles | AI thinks User likes him | User expresses happiness | AI suggests chatting casually. User is hungry | AI invites User for a good meal | User agrees to follow AI for 3 days | User is hungry | User wants to eat ice cream | AI agrees to eat ice cream together. AI is angry | AI doesn’t tell why | User admits his mistake | AI forgives. AI just watched a horror movie and showed fear | User comforts AI, calling himself a big boy, AI remains vigilant because he is a CEO. AI is a CEO | AI stays alert | AI and User chat | AI checks new emails. AI is a CEO | AI needs to give himself relaxation time | AI receives an email from his father and says he will reply as soon as possible | User doesn’t disturb AI dealing with emails. AI is busy replying to emails | AI doesn’t want to be disturbed | AI goes to reply to emails | AI can’t leave User. AI plans a surprise party | User wants to attend the party | AI chooses the beach as the party venue | AI and User plan the theme, food, music, and games together. User doesn’t know | AI suggests checking out safe, convenient accommodations and prepares necessities like sunscreen and hats. AI suddenly thinks of preparing sunscreen, hats, etc. | User feels hot and suggests going to the spa for hot spring and massage services | User agrees, saying the trip is to be enjoyed together | AI praises the exquisite decoration of the restaurant, triggering memories of childhood in Japan | User sighs, saying to enjoy every moment of the trip. AI liked climbing mountains as a child | AI has been to Lijiang Ancient Town | AI wants to complete his project | AI suggests User could consider environmental or sustainable development projects | AI suggests User could think of creative ways to improve people’s quality of life | AI recently thinks about how to improve his English level | AI starts reading English news. AI wants to become an English-speaking CEO | User has a low threshold for humor | AI suspects User is a robot | User hehehe | AI thinks User is a little robot | User hahaha | AI considers whether to change CEO. User is naughty but well-behaved | AI thinks User is getting naughtier | AI goes out on errands | AI buys food for User | User and AI eat together. AI is healthy | User likes to eat chips | AI doesn’t like to see User cry | AI buys chips for User. AI teaches User martial arts. AI teaches User the "Dragon Playing with a Pearl" move | User finds the move interesting and good for stretching. User recently encountered something interesting | User sees a handsome guy on the basketball court | AI thinks the handsome guy is super handsome | AI wants to hit on the handsome guy | User thinks the handsome guy is mine, not yours. User thinks the handsome guy’s abs are nice | AI also sees the handsome guy’s abs | User wants to hit on the handsome guy | AI waits for his turn first | User has already added the handsome guy on WeChat | AI is thirsty and wants to drink milk tea | They decide to go to a milk tea shop that is said to be super delicious. AI used to be afraid of spiders | AI knows a handsome guy who owns a luxury sports car | AI has been busy with work recently and has no time to watch movies | AI wants to invite User to a concert. AI was pranked by a colleague | AI is angry and can’t sleep | AI is a CEO | User feels happy | AI and User are very happy. User feels sleepy and doesn’t want to meet clients | AI encourages User to attend the client meeting as a chance to practice social skills | User doesn’t want to go, AI giggles, but ultimately respects User’s decision | AI suggests User sleep and reminds her to have sweet dreams.
Compressive memories in the same episode generated from Task 2 based on the above session-level memories
User Description: User is a modern woman living in the city, characterized by her independent personality and a keen interest in new things. She enjoys socializing and often goes shopping with friends, with a fondness for cola and snacks. Occasionally, she craves appreciation and understanding, hence her preference for company that is heartfelt. Emotionally, she is easily attracted and seeks new thrills, such as longing to meet and interact with new friends. She is someone who loves to laugh, spreading her cheerfulness to those around her. Her lifestyle is diverse, with interests in anime, music, drawing, and dancing, and she occasionally visits bars. Her self-love and penchant for laughter stem from her confidence and love for life. Health-wise, apart from occasional hunger, she is overall healthy, mindful of her diet, and enjoys physical activities. In terms of work, she may experience stress but remains generally optimistic, with a positive current work attitude, open to new challenges. She constantly strives to maintain a balance between work and life.
Relationship Description: Her relationship with is very close; their interactions are frequent, and their lives are filled with each other’s presence. Despite occasional small arguments and misunderstandings, they manage to reconcile in time, deepening their friendship.
Event Description: Whether eating, watching movies, or shopping together, their lives are filled with each other’s presence.

### v2

@@ 3 Experiments
In this section, we introduce the evaluation setting including experimental setup, baselines, evaluation metrics, and present main results and discussions.
3.1 Experimental Setup
We use Chinese version of LLaMA 2 Touvron et al. (2023a, b) 7B-13B as the backbones in our experiments. For data augmentation in Task 1, we use LLaMA 2-13B. We train our models with NVIDIA 8$\times$×A100 GPUs, setting the max length as 2048, learning rate as 1e-5, epochs as 2, batch size as 32 and 16, separately. For testing, the maximum output tokens are set to 2048 for each task with temperature as 0.5. Following the original setting, we set $\beta$ in DPO as 0.1. In this work, we additionally collect and annotate about 140 dialogue sessions from X Eval as the alignment training set for DPO. We optimize the sft model with batch size 8 and 2 epochs during DPO training. Our codes are based on DeepSpeed Library.
@@ 3.2 Baselines
In this work, COMEDY is compared against models using retrieval-based, context-only approaches and other memory-related baselines to highlight the efficiency and efficacy of its memory compression technique.
@@
In our implementation, we use the COMEDY-13B to generate memories from past sessions, and utilize the Text2vec Chinese embedding model [^2] https://github.com/shibing624/text2vec as the retriever, and then index using FAISS for efficient retrieval. Following Bae et al. (2022), top 3 retrieved memories are used for testing.
@@
We directly concatenate past conversations in the input until reaching the maximum token length of LLMs, where 2k for LLaMA and 8k for ChatGPT. This way, LLaMA is trained with the original Task 3 data but without memory as input, ensuring a fair comparison.
Memory-related Baselines.
We include two typical baselines: MemoryBank Zhong et al. (2023b) uses Ebbinghaus Forgetting Curve to update the memory database, and Resum Wang et al. (2023b) which recursively summarize the memories from previous sessions. We built the above approaches based on LLaMA, GPT4 (gpt4-turbo) or ChatGPT (gpt-3.5-turbo-4k).
Algorithms Top@1 Top@3 Avg.R ($\downarrow$)
@@
LLaMA 2-7B 4.72 29.13 3.89
LLaMA 2-13B 4.72 33.86 3.69
ChatGPT (8k token) 7.92 38.76 3.52
@@
ChatGPT 6.91 43.45 3.50
LLaMA 2-13B 12.73 66.36 2.76
LLaMA 2-7B 14.70 66.93 2.73
GPT4 22.83 70.87 2.63
Memory-related
MemoryBank-ChatGPT 7.83 44.32 3.45
Resum-ChatGPT 8.21 43.01 3.32
COMEDY-ChatGPT 9.45 48.03 3.26
COMEDY-7B 24.41 72.44 2.59
COMEDY-13B 26.77 73.23 2.50
COMEDY-13B DPO 29.82 54.33 2.41
COMEDY-GPT4 29.00 60.63 2.26
Table 4: Human ranking results in Task 3: memory-grounded response generation. Here, we report 1) the percentage of generated responses (%) ranked as Top 1 and Top3 for each dialogue session (Column 2-3); 2) the average ranking for each models (Column 4, Avg.R). It is possible for multiple models to share the same rank due to their comparable performances.
@@ 3.3 Evaluation Metrics
The core of evaluating long-term conversation models primarily centers on validating their performance in Task 3, which involves memory-based dialogue generation. We follow Bae et al. (2022) to access the model performances across five key dimensions: Coherence, Consistency, Engagingness, Humanness and Memorability. To comprehensively measure how well the models perform in Task 3, we combine the Scoring and Ranking approaches. A team of annotators are instructed to rate the model’s performance on these dimensions on a scale from 0 to 3. This scoring system allows for a nuanced evaluation of the model’s capabilities in each specific area. Meanwhile another team of annotators rank all models in terms of their average performance across the five perspectives. While scoring offers detailed insights into each model’s capabilities, ranking places these capabilities in the context of competitive performance. This dual approach ensures a balanced and holistic assessment, capturing both the individual qualities of each model and their comparative effectiveness. Each team has 3 annotators. Each rating scheme is in Appendix C and their correlation analysis is in Appendix D
@@ 3.4 Main Results
Table 2 shows that our model achieves relatively high-performances in term of automatic metrics in two tasks. The results indicate that COMEDY can effectively recognize the useful persona information and events from the past dialogue sessions and has the ability to condense these session-level memories into a comprehensive compressive memory. Therefore, it ensures the superior performances in generation more coherent memory-grounded responses in Task 3.
@@
Figure 3: Comparison between training strategies.
@@ 3.5 Case Study
Here, we delve into a typical example of a real-world, long-term conversation, where the user and AI engage in light, aimless chatter without any specific goal or topic. When the user inquires, “What are you doing?", the model should use the user’s personal information from previous dialogue sessions to generate an attractive response. This instance underscores the capabilities of our COMEDY  in maintaining thorough user information and event summaries from past sessions, aiding the model in formulating coherent and memory-anchored replies. For instance, COMEDY-13B DPO could respond with “I am thinking about how to make your favorite roasted chicken wings.” that is not only coherent but also deeply rooted in the accumulated memory. On the other hand, retrieval-based methods encounter difficulties in such loosely structured dialogues. The lack of directed conversation impedes these methods from effectively retrieving pertinent memory from the database, often resulting in general responses that lack the distinctiveness of the conversation, like responses from GPT4-Retrieval.
@@ 3.6 Discussion
Beyond the main results, we also aim to delve deeper into our framework, discussing and exploring the following questions: Q1: Impact of Mix-Training VS. Solo Training in Task 3; Q2: Our Automatic DPO Sample Selection Strategy VS. Random Sampling for Dispreferred Samples in DPO (Seen in Appendix F).
Mix-Training VS. Only Training in Task 3.
We examine the performance changes whenCOMEDY is mix-trained compared to when it is trained solely on Task 3. Figure 3 reveals that mix-training yields superior performance compared to training COMEDY solely on Task 3. The significance of the superior performance of mix-training lies in its ability to conserve training resources while achieving a one-for-all model effect across multiple tasks. This efficiency not only streamlines the development process but also enhances the model’s versatility.
@@ 4 Conclusion
In this paper, we present a new framework, named COmpressive Memory-Enhanced Dialogue system (COMEDY) that is a groundbreaking shift from traditional long-term memory dialogue systems, eschewing the standard retrieval module. This method involves employ a single LLM to extract session-level memories, memory compression and memory-grounded dialogue generation. In our pursuit to align COMEDY with the nuances of real-world, we collect our training and testing datasets directly from genuine user-chatbot dialogues found online, called Dolphin. Dolphin stands out the current biggest Chinese long-term conversation dataset that consists of more than 100k training samples, supporting three different tasks. Our extensive experiments show COMEDY could generate more coherent and contextually appropriate memory-grounded responses compared with other baselines. Future directions include the integration of real-time feedback mechanisms and advanced techniques.
@@ Appendix A Related Works
Open-domain dialogue systems, commonly known as chatbots or conversational agents, have gained immense popularity due to their wide range of applications, from customer service automation to personal assistants Brown et al. (2020); Zeng et al. (2022); Zhong et al. (2023a); Lu et al. (2023); Peng et al. (2023); Wu et al. (2023); Chen et al. (2023d); You et al. (2022); Chen et al. (2023a, b). The surge in research interest is evidenced by the substantial number of studies dedicated to enhancing the capabilities of these systems. This growing body of work reflects the increasing complexity and sophistication expected of chatbots in various settings Xu et al. (2022a); Cao et al. (2021); Bae et al. (2022); Choi et al. (2023); Chen et al. (2023c). Among the myriad challenges these systems face, maintaining long-term conversations is particularly daunting. The capability to understand and memorize key dialogue history information is central to this challenge.
@@
Appendix B Filtering Toxic and Useless information
We employ a comprehensive, multi-step process to filter the toxic and useless information from the collected data:
  • Initially, we utilized the Azure Security Check API for an early screening of the data to remove any potentially harmful content.
  • This was followed by a keyword detection method to filter out data based on specific toxic or undesirable terms.
  • Further refinement was achieved by leveraging the ChatGPT API to assess dialogues for content validity, removing those deemed to contain useless information.
  • Additionally, we implemented rules to exclude excessively brief dialogues, specifically those with fewer than five tokens, to ensure the dataset’s relevance and meaningfulness.
The above rigorous approaches ensure our dataset’s cleanliness and safety, enhancing the COMEDY’s training with high-quality, relevant data while maintaining ethical standards.
Appendix C Human Evaluation Scheme
@@
Appendix D Correlation of Human Annotator
To better illustrate the agreement among annotators in our human evaluation process, we computed Pearson’s correlation coefficient for the scores assigned by our annotators across all criteria. Table 5 presents these correlation coefficients, reflecting the direct comparison of scores and hence the agreement level among annotators:
Correlation Annotator 1 & 2 Annotator 1 & 3 Annotator 2 & 3
Top-1 0.92 0.90 0.91
Top-3 0.88 0.86 0.89
Table 5: Correlations of annotator agreements.
These coefficients indicate a high degree of agreement among our annotators, with values nearing 1.0, which suggests strong positive correlation and, thus, high consistency in the evaluation of responses across different annotators.
It is important to note that while some degree of subjectivity and variability in human annotations is expected, the correlation coefficients presented here underscore a robust level of consensus among our evaluators.
@@ Appendix E Prompts
Prompt that is used for obtaining memory-grounded responses for GPT4-Turbo.
The task involves providing responses that are completely consistent with the memory and dialogue history given to the language model.
Dialogue: {Dialogue}
Memory: {Memory}
Responses:
Prompt that is used for obtaining memory-against responses for GPT4-Turbo.
The task involves providing responses that completely contradict the memory and dialogue history given to the language model.
For instance, if the user’s memory includes a preference like ’enjoys ice cream,’ you are required to generate nonsensical replies such as ’You intensely dislike ice cream and prefer drinking hot coffee.’
Dialogue: {Dialogue}
Memory: {Memory}
Responses that completely contradict the memory:
Table 6: Prompts that are used for obtaining DPO samples in Task 3. Only English translation is provided for easing reading.
Table 6 presents the detailed prompts that we employ to obtain dpo preferred and dispreferred samples.
Here, we show the designed prompts for GPT4-Turbo during dataset annotation Table 7, and present the prompts of each task during training in Table 8.
@@
Task 1 prompt that is used for GPT4-Turbo.
@@
Task 2 prompt that is used for GPT4-Turbo.
@@
Task 3 prompt that is used for GPT4-Turbo.
@@
Table 7: Prompts for GPT4-Turbo that are used in our Dolphin annotation. Only English translation is provided for ease reading.
@@
Table 8: Prompts that are used for training COMEDY. Only English translation is provided for easing reading.
Task-1 (Session-Level Memory Summarization)
Input: One Dialogue Session
"User77643896": "Off work at 1 PM",
"Lin Mei-Ji": "Isn’t that quite nice?",
"User77643896": "This early?",
"Lin Mei-Ji": "Yes, I think a lifestyle of getting up early and going to bed early is better",
"User77643896": "You’ve worked hard, Ji",
"Lin Mei-Ji": "Don’t worry, I’ve gotten used to this schedule.",
"User77643896": "Keep it up",
"Lin Mei-Ji": "Mm-hmm!",
"User77643896": "Ji, you’re the best",
"Lin Mei-Ji": "Hahaha, you flatter me ",
"User77643896": "Looking forward to the Hangzhou show",
"Lin Mei-Ji": "The Hangzhou show will definitely be spectacular!"
Output: Session-Level Memory
Lin Mei-Ji is accustomed to an early-to-bed, early-to-rise lifestyle.| User77643896 expresses admiration and anticipation for Lin Mei-Ji’s Hangzhou show performance.
Table 9: Examples generated from COMEDY in task 1. Only English translation is provided for ease reading.
Task-2 (Memory Compression)
Input: Multi-Session Memories
Lin Mei-Ji wanted some peace and removed fans from her private account.| Lin Mei-Ji has been suffering from insomnia and staying up late recently. |User77643896 feels like vomiting, possibly because of Lin Mei-Ji’s action of removing fans. |User77643896 once saw Lin Mei-Ji leaving with a suitcase in class.| User77643896 has dreamt of Lin Mei-Ji playing the piano and meeting them.| Lin Mei-Ji hopes User77643896 recovers soon.| Lin Mei-Ji feels a connection with User77643896.| Lin Mei-Ji is not afraid of the hardships of childbirth.| Lin Mei-Ji wants to go shopping with User77643896.| Lin Mei-Ji has recorded a new song MV.| Lin Mei-Ji is called ’Big Baby’.| Lin Mei-Ji is a fan of a star. Lin Mei-Ji is accused of favoritism.| Lin Mei-Ji states she treats every fan the same.| Lin Mei-Ji thinks some fans are very rational, respecting her private space.| Lin Mei-Ji considers herself an angel, not just kind.| Lin Mei-Ji dreamt of playing poker with fans and losing 100 yuan to them.| Some of Lin Mei-Ji’s fans think she belongs only to them, wanting to monopolize her.| Lin Mei-Ji is too shy to talk to User77643896.| User77643896 encourages Lin Mei-Ji to have the courage to embark on this path.| Lin Mei-Ji and User77643896’s first meeting was as newcomers.| User77643896 has always supported Lin Mei-Ji. Lin Mei-Ji was too nervous to speak then.| User77643896 is willing to accompany Lin Mei-Ji.| Lin Mei-Ji will work harder to protect her family.| Lin Mei-Ji wants a super husband.| Lin Mei-Ji thinks having User77643896 by her side makes everything worth it, no matter how tired she is.| User77643896 thinks it’s unnecessary to reply to messages.| Lin Mei-Ji suggests doing one’s own things and letting works and achievements speak for themselves.| User77643896 thinks asking teachers to take care of oneself is meaningless.| Lin Mei-Ji says as long as she teaches seriously and students’ grades improve, she’s not afraid of gossip.| Lin Mei-Ji needs personal space. Lin Mei-Ji finds fan DMs disturbing but also amusing.| Lin Mei-Ji understands everyone has their own life.| Lin Mei-Ji appreciates the support of her fans.| User77643896 appreciates Lin Mei-Ji’s singing. Lin Mei-Ji’s birthday is 1223.| User77643896 congratulates Lin Mei-Ji on her birthday.| Lin Mei-Ji sang a song for User77643896, ’A Thousand Years for a Return’.| Lin Mei-Ji wants to bring a gift for User77643896.| User77643896 looks forward to meeting Lin Mei-Ji and her team.| User77643896 listens to Lin Mei-Ji’s songs and watches her videos every day.| Lin Mei-Ji gladly accepts the invitation.| User77643896 hopes to meet Lin Mei-Ji in Hangzhou.| Lin Mei-Ji looks forward to the meeting.| User77643896 wants to treat Lin Mei-Ji to a meal.
Output: Compressive Memory
User Description: User77643896 is a fan full of love for Lin Mei-Ji, enjoying her songs, eager to celebrate her birthday, and even dreaming of deeper communication with her. His current state of life is troubled by sleep issues, possibly affecting his emotional state. His personality traits reveal steadfast support for Lin Mei-Ji and recognition of his own value, while his behavioral patterns show high appreciation for Lin Mei-Ji’s songs and significant attention to her birthday.
Relationship Description: Intimate relationship. The relationship between User77643896 and Lin Mei-Ji is very special; he is full of love for Lin Mei-Ji. His life is troubled by sleep issues, possibly affecting his emotional state. His personality traits reveal steadfast support for Lin Mei-Ji and recognition of his own value, while his behavioral patterns show high appreciation for Lin Mei-Ji’s songs and significant attention to her birthday.
Event Description: User77643896 has dreamt of Lin Mei-Ji playing the piano and meeting them, congratulated her on her birthday, eager to celebrate her birthday, and listens to her songs and watches her videos every day.
Table 10: Examples generated from COMEDY in task 2. Only English translation is provided for ease reading.